The sum of a number and its reciprocal is given below.
29/10. Find the numbers.

Answer by sudhanshu_kmr(1152)


x + 1/x = 29/10
=> 10x^2 + 10 = 29x
=> 10x^2 -29x +10 =0
=> 10x^2 -25x -4x +10 =0
=> 5x(2x -5) -2(2x-5) = 0
=> (5x -2)(2x-5) = 0
numbers are 2/5 and 5/2.